Determine the volume (in L) of a 0.0764 M solution of K3PO4 that contains 38.7 g of K3PO4.
Answer #1

Molarity = number of moles of K3PO4 / volume of solution in L

number of moles of K3PO4 = 38.7 g / 212.27 g/mol = 0.182 mole

0.0764 = 0.182 / volume of solution in L

volume of solution in L = 0.182 / 0.0764 = 2.38 L

Therefore, the volume of solution = 2.38 L